### v1

@@ 6.3 Issues and future studies
(1) Optimization and trade-offs in performance and computational costs. 3D reconstruction, essential for applications such as measurement, mapping, and monitoring, inherently prioritizes accuracy. Many studies focus on improving algorithmic performance, often at the expense of computational costs. However, as image resolution increases and survey areas expand, processing costs also grow significantly and cannot be overlooked. The ideal objective is to implement a 3D reconstruction solution that can accurately capture scene geometry and texture while utilizing resources efficiently, including runtime, computational load, and equipment costs. For example, MVSFormer++ exemplifies a commendable balance by achieving high accuracy with minimal processing time. In contrast, NeRF-based and GS-based methods, despite their exceptional performance (e.g., GOF), often struggle with significant computational demands, such as high memory usage, extended training times, and limited generalization across different scenes. These challenges constrain their practical application. Besides algorithmic advancements, balancing performance and cost can be achieved through the integration of multiple approaches. For instance, in dense reconstruction of extensive coral reef areas, MVSFormer++ can quickly produce high-quality dense models of the entire area, while methods like GOF can be applied to specific regions of interest to enhance detail and reliability. Additionally, employing strategies such as a coarse-to-fine approach can further optimize the balance between performance and cost.
@@
(2) Evaluation metrics for coral reef 3D reconstruction. Currently, algorithms for evaluating reconstruction quality are underdeveloped, and progress in this area is relatively slow. Coral reefs present a unique and challenging scenario due to their intricate, densely packed structures and varying topography. Additionally, as coral reefs grow slowly—only a few centimeters per year—the reconstruction accuracy needs to achieve millimeter-level precision (Zhong et al. (2023a)). Therefore, accurate assessment of reconstruction results is crucial and necessitates the development of more representative evaluation metrics for quantifying shape reconstruction analysis results (Chen and Zhang (2019)). Future research on evaluation methods should consider a comprehensive range of factors, including global and local perspectives, geometric and radiometric aspects, as well as accuracy and uncertainty. In ecological monitoring, it is crucial to focus on metrics that reflect coral reefs’ ecological functionality, such as biomass, diameter, height, and surface roughness. These indicators can be used to evaluate how well 3D reconstructions capture ecological features, supporting the conservation and management of coral reef ecosystems.
@@
(4) Addressing the effects of light scattering in water. Water inherently causes light scattering, which presents substantial challenges in underwater photography, leading to issues such as color distortion and image blurring. To address this, current approaches typically involve color calibration or advanced image enhancement techniques, including generative adversarial networks (GANs) (Yu et al. (2019)) and diffusion models (Tang et al. (2023)). However, camera color calibration typically targets individual images, making it time-consuming and less effective for multi-view datasets. While image enhancement techniques can produce visually convincing results, they often lack physical accuracy, resulting in inconsistencies between color information and geometric structures. The NeRF framework, however, offers a promising solution to the challenges of light scattering in underwater environments. NeRF methods utilize volumetric rendering, allowing them to model both the geometry and the medium of a scene. For instance, Levy et al. introduced SeaThru-NeRF (Levy et al. (2023)), which integrates a scattering image formation model into the NeRF rendering equations to separate backscatter components from the scene. Similarly, Li et al. developed WaterSplatting (Li et al. (2024)), employing 3D Gaussian Splatting to explicitly represent the scene’s geometry while utilizing a separate volumetric field to capture the water. These methods facilitate the interpretation and modeling of light propagation, scattering, and absorption in underwater environments from a 3D perspective, while optimizing color and geometry. Consequently, they improve consistency between color and geometric information across varying underwater conditions, supporting high-quality 3D reconstruction. Looking ahead, combining these techniques with surface reconstruction methods like SuGaR and GOF could lead to significant advancements in underwater 3D reconstruction.
@@
(5) Demand for datasets and the application of simulators. To advance the study and evaluation of algorithms, it is crucial to have both appropriate metrics and suitable datasets. However, due to the limited distribution of coral reefs and the challenges of underwater photography, datasets for coral reef 3D reconstruction are rare. Even available datasets often lack ground-truth 3D information of the scenes, making it difficult to rigorously assess reconstruction accuracy. This highlights the urgent need for relevant datasets. While obtaining ground-truth data in read-world coral reefs is challenging, simulation offers a viable solution. Recent advance in simulators, particularly in autonomous driving, illustrate this potential (Rong et al. (2020); Hu et al. (2023)). For example, simulators like CARLA enable extensive testing by generating synthetic scenarios, which can also benefit 3D reconstruction (Dosovitskiy et al. (2017)). Synthetic data has proven effective for training models in SuperPoint and R2D2 and for evaluating algorithms in NeRF and Gaussian Splatting research. Therefore, it is entirely reasonable and feasible to design simulation environments and synthetic datasets for coral reef 3D reconstruction, as demonstrated by the datasets generated using AirSim in this study. Future development should consider various factors, including coral morphologies, water quality, and lighting conditions, ensuring the simulator achieves high fidelity to closely mimic real-world environments. Despite advancements, inherent discrepancies remain between simulated and real-world environments, particularly concerning factors such as lighting and texture. Consequently, i research in simulation-to-real transfer and reality gap modeling is essential to effectively address these shortcomings (Daza et al. (2023)).
@@ Acknowledgements
The data for this study was obtained from the Moorea IDEA project. This work was supported by the National Science Fund for Distinguished Young Scholars (grant number 62425102); the U.S. National Science Foundation (grant number OCE 2224354 and earlier awards) for the Moorea Coral Reef LTER; the Gordon and Betty Moore Foundation in support of the MCR LTER; the Italian Minister of University and Research (grant number PNRA18 00263- B2); the National Natural Science Foundation of China (NSFC- 41901407); the Institute of Theoretical Physics, ETH Zurich. We especially would like to thank Prof. Matthias Troyer for his financial and scientific support provided. Special thanks are extended to the ETH Zurich Photogrammetry and Remote Sensing Group and the LIESMARS of Wuhan University.
@@
Name Category Codes SIFT (Lowe (2004)) A hand-crafted local feature detector Refer to the relevant code in the OpenCV library at https://github.com/opencv/opencv KAZE (Alcantarilla et al. (2012b)) A hand-crafted local feature detector Refer to the relevant code in the OpenCV library at https://github.com/opencv/opencv SuperPoint (DeTone et al. (2018)) A deep learning-based local feature detector https://github.com/magicleap/SuperPointPretrainedNetwork R2D2 (Revaud et al. (2019)) A deep learning-based local feature detector https://github.com/naver/r2d2 DISK (Tyszkiewicz et al. (2020)) A deep learning-based local feature detector https://github.com/cvlab-epfl/disk ALIKED (Zhao et al. (2023)) A deep learning-based local feature detector https://github.com/Shiaoming/ALIKED DeDoDe (Edstedt et al. (2024a)) A deep learning-based local feature detector https://github.com/Parskatt/DeDoDe

### v2

@@ 4.3 Issues and future studies
(1) Optimization and trade-offs in performance and computational costs. 3D reconstruction, essential for applications such as measurement, mapping, and monitoring, inherently prioritizes accuracy. Many studies focus on improving algorithmic performance, often at the expense of computational costs. However, as image resolution increases and survey areas expand, processing costs also grow significantly and cannot be overlooked. The ideal objective is to implement a 3D reconstruction solution that can accurately capture scene geometry and texture while utilizing resources efficiently, including runtime, computational load, and equipment costs. For example, MVSFormer++ exemplifies a commendable balance by achieving high accuracy with minimal processing time. In contrast, NeRF-based and GS-based methods, despite their exceptional performance (e.g., GOF), often struggle with significant computational demands, such as high memory usage, extended training times, and limited generalization across different scenes. These challenges constrain their practical application. Besides algorithmic advancements, balancing performance and cost can be achieved through the integration of multiple approaches. For instance, in dense reconstruction of extensive coral reef areas, MVSFormer++ can quickly produce high-quality dense models of the entire area, while methods like GOF can be applied to specific regions of interest to enhance detail and reliability; adopting strategies such as a coarse-to-fine approach can further optimize the balance between performance and cost. Moreover, recent fully feed forward, multi-task geometry predictors such as VGGT (Wang et al. (2025a)), $\pi^{3}$ (Wang et al. (2025b)) and Dense3R (Fang et al. (2025)) accept unposed or unordered image sets and directly infer multiple geometric outputs in a single forward pass without requiring traditional optimization or explicit geometric priors. This unified end-to-end paradigm offers promising robustness, scalability, and efficiency for future underwater 3D reconstruction workflows.
@@
(2) Evaluation metrics for coral reef 3D reconstruction. Currently, algorithms for evaluating reconstruction quality are underdeveloped, and progress in this area is relatively slow. Coral reefs present a unique and challenging scenario due to their intricate, densely packed structures and varying topography. Additionally, as coral reefs grow slowly—only a few centimeters per year—the reconstruction accuracy needs to achieve millimeter-level precision (Zhong et al. (2023a)). Therefore, accurate assessment of reconstruction results is crucial and necessitates the development of more representative evaluation metrics for quantifying shape reconstruction analysis results (Chen and Zhang (2019); Figueira et al. (2015)). Future research on evaluation methods should consider a comprehensive range of factors, including global and local perspectives, geometric and radiometric aspects, as well as accuracy and uncertainty. In ecological monitoring, it is crucial to focus on metrics that reflect coral reefs’ ecological functionality, such as biomass, diameter, height, and surface roughness. These indicators can be used to evaluate how well 3D reconstructions capture ecological features, supporting the conservation and management of coral reef ecosystems.
@@
(4) Addressing the effects of water on imaging. As light propagates through water, it is subject to both absorption and scattering, which significantly degrade image quality. The absorption of light varies across wavelengths, leading to color distortion, while scattering introduces blur and haze (Akkaynak and Treibitz (2019)). In addition, underwater images may be degraded by turbidity and water turbulence, further complicating the imaging process. To mitigate these issues, current approaches typically involve color calibration or advanced image enhancement techniques, including generative adversarial networks (GANs) (Yu et al. (2019)) and diffusion models (Tang et al. (2023)). However, camera color calibration typically targets individual images, making it time-consuming and less effective for multi-view datasets. While image enhancement techniques can produce visually convincing results, they often lack physical accuracy, resulting in inconsistencies between color information and geometric structures. The NeRF framework, however, offers a promising solution to the challenges of light scattering in underwater environments. NeRF methods utilize volumetric rendering, allowing them to model both the geometry and the medium of a scene. For instance, Levy et al. introduced SeaThru-NeRF (Levy et al. (2023)), which integrates a scattering image formation model into the NeRF rendering equations to separate backscatter components from the scene. Similarly, Li et al. developed WaterSplatting (Li et al. (2024)), employing 3D Gaussian Splatting to explicitly represent the scene’s geometry while utilizing a separate volumetric field to capture the water. These methods facilitate the interpretation and modeling of light propagation, scattering, and absorption in underwater environments from a 3D perspective, while optimizing color and geometry. Consequently, they improve consistency between color and geometric information across varying underwater conditions, supporting high-quality 3D reconstruction. Looking ahead, combining these techniques with surface reconstruction methods like SuGaR and GOF could lead to significant advancements in underwater 3D reconstruction.
(5) Multi-source data fusion for underwater 3D reconstruction. In addition to photogrammetry, other sensing approaches such as sonar and LiDAR have been utilized for underwater 3D reconstruction. Sonar enables rapid, large-scale surveys and is useful for broad-area mapping, though its resolution is insufficient for capturing fine-scale structures when compared to image-based methods, which can achieve millimeter-level accuracy. Airborne LiDAR has been widely used in coastal and nearshore ocean mapping and can recover both the ocean surface and seabed topography (Massot-Campos and Oliver-Codina (2015)). Nevertheless, it is limited by relatively low resolution and is susceptible to water surface refraction. In contrast, underwater LiDAR can operate in close proximity to the scene, reducing water-related distortion and allowing for direct distance measurement (Yang et al. (2022); McLeod et al. (2013)). This makes underwater LiDAR particularly valuable as a complement to image-based approaches, as its depth information aids in modeling light propagation through water. However, underwater LiDAR remains relatively underdeveloped and is mainly used as a supplementary tool. With further advancements, particularly toward performance levels comparable to terrestrial LiDAR, it holds great promise for improving the reconstruction of complex structures such as coral reefs, offering enhanced accuracy and robustness through data fusion.
@@
(6) Demand for datasets and the application of simulators. To advance the study and evaluation of algorithms, it is crucial to have both appropriate metrics and suitable datasets. However, due to the limited distribution of coral reefs and the challenges of underwater photography, datasets for coral reef 3D reconstruction are rare. Even available datasets often lack ground-truth 3D information of the scenes, making it difficult to rigorously assess reconstruction accuracy. This highlights the urgent need for relevant datasets. While obtaining ground-truth data in real-world coral reefs is challenging, simulation offers a viable solution. Recent advance in simulators, particularly in autonomous driving, illustrate this potential (Rong et al. (2020); Hu et al. (2023)). For example, simulators like CARLA enable extensive testing by generating synthetic scenarios, which can also benefit 3D reconstruction (Dosovitskiy et al. (2017)). Synthetic data has proven effective for training models in SuperPoint and for evaluating algorithms in NeRF and Gaussian Splatting. Therefore, it is both reasonable and feasible to develop simulation environments and synthetic datasets for coral reef 3D reconstruction, as demonstrated by the datasets generated using AirSim in this study. Several works have contributed to this area. GEODT (Nakath et al. (2022)), an optical digital twin for underwater environments, supports the development and validation of refractive photogrammetric algorithms. POSER (Menna et al. (2024)) was proposed for teaching and training underwater photogrammetry, allowing users to explore various acquisition strategies. Potokar et al. proposed HoloOcean (Potokar et al. (2022)), an underwater robotics simulator based on Unreal Engine 4 that supports the generation of photorealistic underwater imagery. Amer et al. introduced Unav-Sim (Amer et al. (2023)), which enables the simulation of different underwater scenarios and models. In March 2025, Song et al. presented OceanSim (Song et al. (2025)), a high-fidelity simulator designed to improve physics-based underwater sensor modeling. Future development should consider various factors, including coral morphologies, water quality, and lighting conditions, ensuring the simulator achieves high fidelity to closely mimic real-world environments. Despite advancements, inherent discrepancies remain between simulated and real-world environments, particularly concerning factors such as lighting and texture. Consequently, research in simulation-to-real transfer and reality gap modeling is essential to effectively address these shortcomings (Daza et al. (2023)).
@@ Acknowledgements
The data for this study was obtained from the Moorea IDEA project. This work was supported by the National Science Fund for Distinguished Young Scholars (grant number 62425102); the U.S. National Science Foundation (grant number OCE 2224354 and earlier awards) for the Moorea Coral Reef LTER; the Gordon and Betty Moore Foundation in support of the MCR LTER; the Italian Minister of University and Research (grant number PNRA18 00263- B2); the National Natural Science Foundation of China (NSFC- 41901407); the Institute of Theoretical Physics, ETH Zurich. We especially would like to thank Prof. Matthias Troyer for his financial and scientific support provided. Special thanks are extended to the ETH Zurich Photogrammetry and Remote Sensing Group. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
@@
Name Category Codes SIFT (Lowe (2004)) A hand-crafted local feature detector Refer to the relevant code in the OpenCV library at https://github.com/opencv/opencv KAZE (Alcantarilla et al. (2012)) A hand-crafted local feature detector Refer to the relevant code in the OpenCV library at https://github.com/opencv/opencv SuperPoint (DeTone et al. (2018)) A deep learning-based local feature detector https://github.com/magicleap/SuperPointPretrainedNetwork R2D2 (Revaud et al. (2019)) A deep learning-based local feature detector https://github.com/naver/r2d2 DISK (Tyszkiewicz et al. (2020)) A deep learning-based local feature detector https://github.com/cvlab-epfl/disk ALIKED (Zhao et al. (2023)) A deep learning-based local feature detector https://github.com/Shiaoming/ALIKED DeDoDe (Edstedt et al. (2024a)) A deep learning-based local feature detector https://github.com/Parskatt/DeDoDe